# GSplatVNM: Point-of-View Synthesis for Visual Navigation Models Using Gaussian Splatting

### v1

@@ II-C Visual Navigation with Neural Rendering Models
To overcome the limitations of ITGs, a promising direction is to use neural rendering models as a prior map for visual navigation. Some studies generate target point-of-view images online using world models [19, 20], but these methods require significant computation and lack geometric information about the environment. In contrast, Neural Radiance Fields (NeRF) [21] and 3DGS [22] provide geometric information and can quickly generate novel point-of-view images, making them well-suited for navigation tasks. In particular, 3DGS relies on lightweight 3D Gaussians as its physical representation, which have been used as a geometric prior map for visual navigation [22, 23, 24, 25]. However, these studies assume online self-localization on 3DGS. In contrast, this study integrates 3DGS and VNM to use a renderable and geometric environment representation as an alternative to ITG, while benefiting from VNM to eliminate the need for online self-localization.
@@ III-A Representation by 3D Gaussians
3DGS is an advanced technique for the reconstruction and rendering of environments. Instead of using traditional mesh or voxel-based approaches, environments are encoded as a set of 3D Gaussians, each associated with color, opacity, and covariance parameters. An environment is represented by 3D Gaussians, denoted as $\{(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{c}_{i},\alpha_{i})\}_{i=1}^{N}$, where $N$ is the number of Gaussians, $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$ is the mean, $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix, $\mathbf{c}_{i}\in\mathbb{R}^{3}$ is the color, and $\alpha_{i}\in[0,1]$ is the opacity of the $i$-th Gaussian. Each Gaussian can be expressed as a continuous function over the 3D position space $\mathbf{x}\in\mathbb{R}^{3}$ as follows: $G_{i}(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})\right).$ To render a view from a given camera pose, we splat the Gaussians onto the 2D image plane and determine the pixel colors using $\mathbf{c}_{i}$ and $\alpha_{i}$. By training the parameters of 3DGS to minimize a reprojection error across multiple views, we obtain a compact and renderable representation of the environment.
@@ III-B Distance Estimation between Robot and 3DGS
3DGS can not only render novel views but also be employed to represent the geometry of the environment [22, 24, 23]. In this work, following the approach proposed in [24], we convert the 3DGS into a set of ellipsoids $\mathcal{E}_{i}$ to estimate the distance between the robot and the environment, where each ellipsoid $\mathcal{E}_{i}$ is defined as $\mathcal{E}_{i}=\{\mathbf{x}\in\mathbb{R}^{3}\mid(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})\leq 1\}.$ Thus, the sphere-to-ellipsoid distance between the $i$-th ellipsoid and the robot centroid $\hat{\mathbf{p}}=\textbf{{F}}\textbf{{R}}^{\top}(\mathbf{p}-\mathbf{\mu}_{i})$ can be formulated as the following quadratic program:
$d_{i}(\mathbf{x}_{r})=\min_{\mathbf{x}\in\mathbb{R}^{3}}\|\mathbf{x}-\hat{\mathbf{p}}\|_{2}^{2},\quad\text{s.t.}\;\mathbf{x}^{\top}\textbf{{S}}^{-2}\mathbf{x}\leq 1,$ (1)
where $\mathbf{x}_{r}$ is the spherical-shaped robot position, F is a reflection matrix (a diagonal matrix with entries of either 1 or $-1$ used to reflect points across the primary planes), R is the rotation matrix of the ellipsoid, and S is a scaling matrix determined by the confidence of the Gaussian, which serves as a hyperparameter.
@@ IV VISUAL NAVIGATION WITH 3DGS
Figure 2: Overview of the proposed GSplatVNM. In a conventional ITG-based approach, the environment is represented by ITG, and the target point-of-views given to the VNM (NoMaD [3]) are limited to a set of the collected image DB. On the other hand, GSplatVNM represents the environment as 3DGS to synthesize dense and traversable point-of-views from the start to the goal, thereby guiding NoMaD. As a result, GSplatVNM can achieve efficient image-goal navigation by seamlessly bridging gaps in the image DB, even when the image DB is sparse.
@@ IV-A Overview of the Proposed Framework
Figure 2 shows an overview of our proposed navigation framework integrating 3DGS with a VNM. Given the start and goal images, we first estimate the robot’s start and goal poses in 3DGS and then plan a global trajectory within the 3DGS representation. We subsequently render target point-of-view images along the planned trajectory to condition the VNM, i.e., NoMaD [3], which is a state-of-the-art VNM. During navigation, NoMaD generates spatial waypoints from the observation images and the synthesized target point-of-view images, and the robot follows these waypoints using MPPI [26].
@@ IV-B Start and Goal Pose Estimation in 3DGS
In the proposed framework, we first estimate the start and goal poses in 3DGS from the given start and goal images. We assume that the robot moves on a 2D plane; hence, the robot pose is represented as $\mathbf{p}=(x,y,\theta_{\rm{yaw}})\in\mathrm{SE}(2),$ where $x$ and $y$ denote the 2D position and $\theta_{\rm{yaw}}$ the yaw angle. Although some studies have addressed accurate and online pose estimation in 3DGS [22], we require only a one-shot, rough estimation of the global pose.
In this work, we optimize the global start and goal poses, $\mathbf{p}_{\rm{start}}$ and $\mathbf{p}_{\rm{goal}}$, by minimizing the following loss function $\mathcal{L}$:
$\mathcal{L}(\mathbf{p}_{*})=\mathcal{L}_{\text{img}}(I_{*},I_{\text{rendered}}(\mathbf{p}_{*}))+\mathbf{1}_{\text{collision}}[d_{\text{min}}(\mathbf{p}_{*})\leq r],$ (2)
$d_{\text{min}}(\mathbf{p}_{*})=\min_{i\in\{0,\dots,N-1\}}d_{i}(\mathbf{p}_{*}),\quad*\in\{\rm{start},\rm{goal}\},$
where $I_{*}$ is the given start or goal image, $I_{\text{rendered}}(\mathbf{p}_{*})$ is the image rendered from 3DGS at pose $\mathbf{p}_{*}$, and $r$ is the radius of the robot. To render the image, in addition to the robot’s 2D pose $\mathbf{p}_{*}$, the $z$ position $z_{*}$, roll $\theta^{*}_{\rm{roll}}$, and pitch $\theta^{*}_{\rm{pitch}}$ angles are required. We assume that the image DB was captured with fixed $(z,\theta_{\rm{roll}},\theta_{\rm{pitch}})$ values, and we use the mean of those values obtained from the estimated camera poses during the training of 3DGS.
The first term, $\mathcal{L}_{\text{img}}$, in (2) is an image similarity metric between the given and rendered images from 3DGS at the pose $\mathbf{p}_{*}$. Specifically, we use the Learned Perceptual Image Patch Similarity (LPIPS) metric [27], which is computed from the feature maps of AlexNet [28] and ranges from 0 to 1. The second term is a collision penalty to avoid the infeasibility of global planning. Here, $\mathbf{1}_{\text{collision}}$ denotes an indicator function, and $d_{i}$ is the distance between the robot and the $i$-th ellipsoid-approximated 3DGS defined in (1).
Although the loss function (2) is theoretically differentiable, its complexity may lead to convergence at local optima. Given that the optimization is over a six-dimensional space, we employ a black-box optimization method known as the Tree-structured Parzen Estimator [29]. The entire optimization process is implemented using Optuna [30].
@@ IV-C Global Trajectory Planning in 3DGS
We then plan a traversable global trajectory in 3DGS from the estimated start to goal poses, $\mathbf{p}_{\rm{start}}$ and $\mathbf{p}_{\rm{goal}}$. In this work, considering both the guarantee of optimality and ease of implementation, we plan a global path $\{(x_{\rm{start}},y_{\rm{start}}),(x^{0},y^{0}),\dots,(x_{M-1},y_{M-1}),(x_{\rm{goal}},y_{\rm{goal}})\}$ using A* search [31] without considering yaw angles, and then assign yaw angles along the path. That is, we compute the traversable trajectory $\mathcal{P}$ as follows:
$\mathcal{P}=\{\mathbf{p}_{\rm{start}},(x_{0},y_{0},\theta_{\rm{yaw}}^{0}),\dots,(x_{M-1},y_{M-1},\theta_{\rm{yaw}}^{M-1}),\mathbf{p}_{\rm{goal}}\},$ (3)
$\{(x_{0},y_{0}),\dots,(x_{M-1},y_{M-1})\}=\text{A*}(\mathbf{p}_{\rm{start}},\mathbf{p}_{\rm{goal}}),$
@@
After the global planning, we render the target point-of-views $\mathcal{{V}}_{r}=\{I_{0}^{r}(\mathbf{p}_{0}),\dots,I_{M-1}^{r}(\mathbf{p}_{M-1}),I_{\rm{goal}}\}$ along the planned trajectory $\mathcal{P}$ using 3DGS. These rendered images are then used to condition the NoMaD policy to generate spatial waypoints.
@@ IV-D Zero-shot Local Planning and Control with NoMaD
NoMaD [3] is a visual goal-conditioned policy that generates spatial waypoints from a sequence of observation images at time $t$, $\mathcal{O}_{\rm{obs}}=\{I_{t-p},\dots,I_{t}\}$ (with $p=3$ in NoMaD), as well as from the target point-of-view image $I_{\rm{goal}}$. NoMaD consists of three networks:
  • A ViNT encoder [2], $\mathbf{c}_{t}=f_{\rm{enc}}(\mathcal{O}_{\rm{obs}},I_{\rm{goal}})$, that extracts goal-conditioned context features from the observation and target images.
  • A diffusion model [33]-based policy, $\pi(\mathbf{c}_{t})$, that generates a set of waypoints $\mathcal{W}_{t}=\{\mathbf{w}_{0},\dots,\mathbf{w}_{L-1}\}$ (with $L=8$ in NoMaD) from the encoded features, where each waypoint represents a 2D physical-scaled position in the robot coordinate.
@@
These networks are trained in a supervised manner on large, heterogeneous datasets collected across a diverse set of environments and robotic platforms over 100 hours, resulting in zero-shot navigation capability on unseen environments and robots.
During navigation, after synthesizing the target point-of-view images, we estimate the nearest image in the rendered target point-of-views, $I_{t}^{r}\in\mathcal{V}_{r}$, based on the current observation image $\mathcal{O}_{\rm{obs}}$ and the previous target image $I_{t-1}^{r}$ (with $I_{0}^{r}=I_{\rm{start}}$). Then, we generate the next point-of-view-conditioned waypoints $\mathcal{W}_{t}$ using the next target image $I_{t+1}^{r}$ according to the policy $\pi(\mathbf{c}_{t})$, where $\mathbf{c}_{t}=f_{\rm{enc}}(\mathcal{O}_{\rm{obs}},I_{t+1}^{r})$. The robot follows these waypoints using MPPI [26], a sampling-based model predictive control method that minimizes future deviations from the waypoints while taking into account the robot’s differential two-wheel dynamics.
@@ V-B2 Pre-Collection of Image DB
For the image-goal navigation task, the robot utilizes a pre-collected image DB as prior knowledge of the environment. To evaluate robustness with respect to the size of the image DB, we prepare multiple image DBs for each environment. Specifically, we collect $\{300,500,1000\}$ images for Greigsville, $\{500,800,1000\}$ images for Ribera, and $\{300,1000,2000,3000\}$ images for skokloster-castle. Prior to the navigation task, the robot explores the environment using a non-goal-conditioned policy based on NoMaD and collects images at regular intervals of 0.5 s until these numbers of images are collected.
@@ VI CONCLUSION
Despite these benefits, several challenges remain. One challenge is improving the tracking performance of the generated point-of-view images in VNM. Approaches to address this issue include fine-tuning NoMaD itself [2] and re-planning the global trajectory [37]. Another challenge is handling dynamic obstacles. Although the current VNM can avoid dynamic obstacles, online reconstruction of 3DGS for dynamic environments [38] could provide a more robust solution. Finally, enhancing the global position estimation accuracy in 3DGS is important. Since image similarity-based methods can sometimes confuse visually similar locations, incorporating instance segmentation [23, 39] may help overcome this limitation.

### v3

@@ II-C Visual Navigation with Neural Rendering Models
To overcome the limitations of ITGs, a promising direction is to use neural rendering models as a prior map for visual navigation. Some studies generate target point-of-view images online using world models [19, 20], but these methods require significant computation and lack geometric information about the environment. In contrast, Neural Radiance Fields (NeRF) [21] and 3DGS [22] provide geometric information and can quickly generate novel point-of-view images, making them well-suited for navigation tasks. In particular, 3DGS relies on lightweight 3D Gaussians as its physical representation, which have been used as a geometric prior map for visual navigation [22, 23, 24, 25].
While such recent studies [22, 23, 24, 25] also use 3DGS for navigation, their approaches typically require online and presize self-localization within the 3DGS map. In contrast, our method uses 3DGS as an offline environment model solely to synthesize a sequence of target viewpoints. A powerful, pre-trained VNM then follows these images without needing to explicitly track its pose against the map. Our core contribution is therefore the integration of 3DGS as a viewpoint generator to guide a localization-free policy, rather than using it as a map for online localization.
@@ III-A Representation by 3D Gaussians
3DGS is an advanced technique for the reconstruction and rendering of environments. Instead of using traditional mesh or voxel-based approaches, environments are encoded as a set of 3D Gaussians, each associated with color, opacity, and covariance parameters. An environment is represented by 3D Gaussians, denoted as $\{(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{c}_{i},\alpha_{i})\}_{i=1}^{N}$, where $N$ is the number of Gaussians, $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$ is the mean, $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix, $\mathbf{c}_{i}\in\mathbb{R}^{3}$ is the color, and $\alpha_{i}\in[0,1]$ is the opacity of the $i$-th Gaussian. Each Gaussian can be expressed as a continuous function over the 3D position space $\mathbf{x}\in\mathbb{R}^{3}$ as follows: $G_{i}(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{\top}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})\right).$ To render a view from a given camera pose, we splat the Gaussians onto the 2D image plane and determine the pixel colors using $\mathbf{c}_{i}$ and $\alpha_{i}$. By training the parameters of 3DGS to minimize a reprojection error across multiple views, we obtain a compact and renderable representation of the environment.
@@ III-B Distance Estimation between Robot and 3DGS
3DGS can not only render novel views but also be employed to represent the geometry of the environment [22, 24, 23]. In this work, following the approach proposed in [24], we convert the 3DGS into a set of ellipsoids $\mathcal{E}_{i}$ to estimate the distance between the robot and the environment, where each ellipsoid $\mathcal{E}_{i}$ is defined as $\mathcal{E}_{i}=\{\mathbf{x}\in\mathbb{R}^{3}\mid(\mathbf{x}-\mathbf{\mu}_{i})^{\top}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})\leq 1\}.$ Thus, the ellipsoid-to-sphere distance between the $i$-th ellipsoid and the robot centroid $\hat{\mathbf{p}}=\textbf{{F}}\textbf{{R}}^{\top}(\mathbf{p}-\mathbf{\mu}_{i})$ can be formulated as the following quadratic program:
$d_{i}(\mathbf{p})=\min_{\mathbf{x}\in\mathbb{R}^{3}}\|\mathbf{x}-\hat{\mathbf{p}}\|^{2},\quad\text{s.t.}\;\mathbf{x}^{\top}\textbf{{S}}^{-2}\mathbf{x}\leq 1,$ (1)
where $\mathbf{p}\in\mathbb{R}^{3}$ is the robot 3D position approximated by a sphere, F is a reflection matrix (a diagonal matrix with entries of either 1 or 1-1- 1 used to reflect points across the primary planes), R is the rotation matrix of the ellipsoid, and S is a scaling matrix determined by the confidence of the Gaussian, which serves as a hyperparameter.
@@ IV VISUAL NAVIGATION WITH 3DGS
Figure 2: Overview of the proposed GSplatVNM. In a conventional ITG-based approach, the environment is represented by ITG, and the target point-of-views given to the VNM (NoMaD [3]) are limited to a set of collected image DB. On the other hand, GSplatVNM represents the environment as 3DGS to synthesize dense and traversable point-of-view images from the start to the goal, thereby guiding NoMaD. As a result, GSplatVNM can achieve efficient image-goal navigation by seamlessly bridging gaps in the image DB, even when the image DB is sparse.
@@ IV-A Overview of the Proposed Framework
Figure 2 shows an overview of our proposed navigation framework integrating 3DGS with a VNM. Given the start and goal images, we first estimate the robot’s start and goal poses in 3DGS and then plan a global trajectory within the 3DGS representation. We subsequently render target point-of-view images along the planned trajectory to condition the VNM, i.e., NoMaD [3], which is a state-of-the-art VNM. During navigation, NoMaD generates spatial waypoints from the observation images and the synthesized target point-of-view images, and the robot follows these waypoints using Model Predictive Path Integral control (MPPI) [26].
@@ IV-B Start and Goal Pose Estimation in 3DGS
In the proposed framework, we first estimate the start and goal poses in 3DGS from the given start and goal images. We assume that the robot moves on a 2D plane; hence, the robot’s pose is represented as $\mathbf{q}=(x,y,\theta_{\rm{yaw}})\in\mathrm{SE}(2),$ where $x$ and $y$ denote the 2D position and $\theta_{\rm{yaw}}$ the yaw angle. Although some studies have addressed accurate and online pose estimation in 3DGS [22], we require only a one-shot, rough estimation of the global pose.
In this work, we optimize the global start and goal poses, $\mathbf{q}_{\rm{start}}$ and $\mathbf{q}_{\rm{goal}}$, by minimizing the following loss function $\mathcal{L}$:
$\mathcal{L}(\mathbf{q}_{*})=\mathcal{L}_{\text{img}}(I_{*},I_{\text{rendered}}(\mathbf{q}_{*}))+\mathbf{1}_{\text{collision}}[d_{\text{min}}(\mathbf{q}_{*})\leq r],$ (2)
$d_{\text{min}}(\mathbf{q}_{*})=\min_{i\in\{0,\dots,N-1\}}d_{i}(\mathbf{q}_{*}),\quad*\in\{\rm{start},\rm{goal}\},$
where $I_{*}$ is the given start or goal image, $I_{\text{rendered}}(\mathbf{q}_{*})$ is the image rendered from 3DGS at pose $\mathbf{q}_{*}$, and $r$ is the radius of the robot. The first term, $\mathcal{L}_{\text{img}}$, in (2) is an image similarity metric between the given and rendered images from 3DGS at the pose $\mathbf{q}_{*}$. Specifically, we use the Learned Perceptual Image Patch Similarity (LPIPS) metric [27], which is computed from the feature maps of AlexNet [28] and ranges from 0 to 1. The second term is a collision penalty to avoid the infeasibility of global planning. Here, $\mathbf{1}_{\text{collision}}$ denotes an indicator function, and $d_{i}$ is the distance between the robot and the $i$-th ellipsoid-approximated 3DGS defined in (1).
To render the image and calculate the distance, in addition to the robot’s 2D pose $\mathbf{q}_{*}$, the $z$ position $z_{*}$, roll $\theta^{*}_{\rm{roll}}$, and pitch $\theta^{*}_{\rm{pitch}}$ angles are required. We assume that the image DB was captured with fixed $(z,\theta_{\rm{roll}},\theta_{\rm{pitch}})$ values, and we use the mean of those values obtained from the estimated camera poses during the training of 3DGS.
Although the loss function (2) is theoretically differentiable, its complexity may lead to convergence at local optima. Given that the optimization is over a three-dimensional space, we employ a black-box optimization method known as the Tree-structured Parzen Estimator [29]. The entire optimization process is implemented using Optuna [30].
@@ IV-C Global Trajectory Planning in 3DGS
We then plan a traversable global trajectory in 3DGS from the estimated start to goal poses, $\mathbf{q}_{\rm{start}}$ and $\mathbf{q}_{\rm{goal}}$. In this work, considering both the guarantee of optimality and ease of implementation, we plan a global path $\{(x_{\rm{start}},y_{\rm{start}}),(x^{0},y^{0}),\dots,(x_{M-1},y_{M-1}),(x_{\rm{goal}},y_{\rm{goal}})\}$ using A* search [31] without considering yaw angles, and then assigning yaw angles along the path. That is, we compute the traversable trajectory $\mathcal{T}$ as follows:
$\mathcal{T}=\{\mathbf{q}_{\rm{start}},(x_{0},y_{0},\theta_{\rm{yaw}}^{0}),\dots,(x_{M-1},y_{M-1},\theta_{\rm{yaw}}^{M-1}),\mathbf{q}_{\rm{goal}}\},$ (3)
$\{(x_{0},y_{0}),\dots,(x_{M-1},y_{M-1})\}=\text{A*}(\mathbf{q}_{\rm{start}},\mathbf{q}_{\rm{goal}}),$
@@
After the global planning, we render a set of the target point-of-views $\mathcal{{V}}_{r}=\{I_{0}^{r}(\mathbf{q}_{0}),\dots,I_{M-1}^{r}(\mathbf{q}_{M-1}),I_{\rm{goal}}\}$ along the planned trajectory $\mathcal{T}$ using 3DGS. These rendered images are then used to condition the NoMaD policy to generate spatial waypoints.
@@ IV-D Zero-shot Local Planning and Control with NoMaD
NoMaD [3] is a visual subgoal-conditioned policy that generates spatial waypoints from a sequence of observation images at time (t), $\mathcal{O}_{\rm{obs}}=\{I_{t-p},\dots,I_{t}\}$ (with p=3 in NoMaD), as well as from the target subgoal point-of-view image $I_{\rm{target}}$. NoMaD consists of three networks:
  • A subgoal image-conditioned vision encoder, $\mathbf{c}_{t}=f_{\rm{enc}}(\mathcal{O}_{\rm{obs}},I_{\rm{target}})$, that extracts context features from the observation $\mathcal{O}_{\rm{obs}}$ and target subgoal image $I_{\rm{target}}$.
  • A diffusion model [33]-based policy, $\pi(\mathbf{c}_{t})$, that generates a set of waypoints $\mathcal{W}_{t}=\{\mathbf{w}_{0},\dots,\mathbf{w}_{L-1}\}$ (with L=8 in NoMaD) from the encoded features $\mathbf{c}_{t}$, where each waypoint represents a 2D physically-scaled position in the robot’s coordinate frame.
@@
These networks are trained in a supervised manner on large, heterogeneous datasets collected across a diverse set of environments and robotic platforms over 100 hours, resulting in zero-shot navigation capability in unseen environments and on unseen robot setups [3].
In our navigation scheme, after synthesizing a set of target point-of-view images toward the goal, $\mathcal{{V}}_{r}$, we first estimate the nearest image $I_{t}^{r}\in\mathcal{V}_{r}$ to the current observation. This estimation is performed using the distance-estimation network, $\phi$, which is conditioned on both the current observation $\mathcal{O}_{\rm{obs}}$ and the target images $I_{i}^{r}\in\mathcal{{V}}_{r}$ (with $I_{0}^{r}=I_{\rm{start}}$). Then, using the subsequent image in the sequence, $I_{t+1}^{r}$, as the next subgoal, we generate point-of-view-conditioned waypoints $\mathcal{W}_{t}$ according to the policy $\pi(\mathbf{c}_{t})$, where $\mathbf{c}_{t}=f_{\rm{enc}}(\mathcal{O}_{\rm{obs}},I_{t+1}^{r})$. The robot follows these waypoints using MPPI [26], a sampling-based model predictive control method that minimizes future deviations from the waypoints while taking into account the robot’s differential two-wheel dynamics.
@@ V-B2 Pre-Collection of Image DB
[…]
@@ V-C2 Baseline Methods
Our goal is to specifically assess the impact of the underlying environment representation (3DGS vs. ITG) on a state-of-the-art zero-shot navigation policy, rather than conducting a broad comparison of different navigation policies such as GNM [1] or ViNT [2]. Therefore, we utilize variants of the same NoMaD policy for all methods to ensure a fair comparison of the representation itself.
@@ VI CONCLUSION
Despite these promising results, our work has several limitations:
  • Real-World Validation: The current validation is confined to simulation environments. A crucial next step is to conduct experiments on a physical robot to evaluate the sim-to-real transferability and robustness of our framework under real-world conditions.
  • Planar Motion Assumption: Our framework assumes the robot operates on a 2D plane, making it unsuitable for environments with significant elevation changes, such as ramps or multiple floors. While our approach could theoretically be extended by incorporating 3D pose estimation and global planning, we leave this extension for future work.
  • Collision Checking with 3DGS: The collision checking in our method lacks geometric precision. 3DGS is optimized for rendering quality rather than metrically accurate reconstruction, which can cause a scale discrepancy between the model and the physical world. Consequently, the robot’s radius for collision checking within the 3DGS space is treated as a hyperparameter in our experiment. However, this limitation is mitigated by the fact that the VNM policy follows rendered target images instead of precise world coordinates, reducing the criticality of exact geometric collision checking.
Looking forward, several technical challenges remain to be addressed. One is improving the tracking performance of the VNM on the generated point-of-view images. Potential approaches include fine-tuning the VNM itself [2] or re-planning the global trajectory [37]. Another challenge is handling dynamic obstacles. Although the current VNM can avoid obstacles, online reconstruction of 3DGS for dynamic environments [38] could provide a more robust solution. Finally, enhancing the global position estimation accuracy within the 3DGS representation is important. Since image similarity-based methods can sometimes confuse visually similar locations, incorporating instance segmentation [23, 39] may help overcome this limitation.